A closely wound 0.20m × .20 m square coil of 100 turns is exposed to an external magnetic field of 0.74 T that is perpendicular to the plane of the coil. The coil is then rotated for 10 s until it makes 60 degrees with the magnetic field. Find the average induced electromotive force during the 10 s interval.

1
Expert's answer
2022-05-20T04:53:56-0400
"{\cal E}=-\frac{\Delta \Phi}{\Delta t}=-\frac{NBA(\cos\theta_2-\cos\theta_1)}{\Delta t}"

"=-\frac{100*0.74*0.20^2(\cos30^\circ-\cos0^\circ)}{10}=0.04\:\rm V"